OpenStudy (anonymous):

The sum of a first and second number is 130. If the second number is 20 less than four times the first number, find the two numbers.

OpenStudy (nikato):

So first set up 2 equations Let a= first number and b= second number We know a+b=130 b= 4a-20 Do you know how to solve from here?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

@nikato

OpenStudy (nikato):

Have you learned substitution?

OpenStudy (anonymous):

yea but can you solve it @nikato

OpenStudy (nikato):

So using substitution, a+4a-20=130 +20. +20 ------------------ a+4a=150. Combine like terms 5a=150. Divide both sides by 5 So a=?

OpenStudy (anonymous):

a=30 @nikato

OpenStudy (nikato):

Sorry. This site crashed on me. I couldn't even message you. @cutegirl And yes a=30 To find b, substitute this value of a into one of the original equation Lets use a+b=130 So 30+b=130 So b=100